OpenStudy (anonymous):

maximum value of sin^2(120+Q)+sin^2(120-Q)

OpenStudy (anonymous):

uhhhh what math is this!!???

OpenStudy (anonymous):

trigonometry

OpenStudy (anonymous):

oh dang i havnt taken this!

OpenStudy (anonymous):

trigonometry means too confusion.

OpenStudy (anonymous):

not an issue guys.. i will try to solve it. Thanks

OpenStudy (anonymous):

i am sorry @sachin

myininaya (myininaya):

Have you tried rewriting it first in some other way

myininaya (myininaya):

I would probably start by using sum and difference identities for sin then square the result then see some things cancel then write in terms of one trig function

OpenStudy (tylerd):

it cant exceed 2

OpenStudy (tylerd):

have a hunch that when Q = 0 thats the max.

myininaya (myininaya):

@tylerd for some reason that would work but I don't know why.

OpenStudy (tylerd):

same

myininaya (myininaya):

I found the same answer doing the whole rewriting thing

myininaya (myininaya):

So it was just a hunch @TylerD Any clue why that works?

OpenStudy (tylerd):

plugged in some values and noticed the closer it got to q=0, the higher the value. if you go backward it gets smaller then eventually comes full circle. or something like that.

myininaya (myininaya):

Ah. Limits.

OpenStudy (tylerd):

but i suppose q could also equal 360 or -360?

myininaya (myininaya):

well I know it is I won't show exactly how I rewrote it But I will tell you I wrote sin some form where you have B-Asin^2(theta) where A and B are some constants I would like @sachin.davra to obtain himself Then I thought of what would be sin^2's min since we are subtracting there And if for some reason you had found A to be negative then you would actually find the sin^2's max. Then add. Also look that some expression and had fun with some calculus giving me the same result

myininaya (myininaya):

hint: both A and B >0

myininaya (myininaya):

sin's max is 1 sin^2's max is also 1

OpenStudy (tylerd):

may have something to do with sin^2x+cos^2x=1 where sin(90-x)=cos(x)

myininaya (myininaya):

sin^2's min is 0

myininaya (myininaya):

But sachin.davra posted this like 18 hours ago he is probably no longer here

OpenStudy (tylerd):

oh wow, says hes here though maybe afk.

OpenStudy (dawnr):

i suck at it too so if anyone could help as well that would be great http://openstudy.com/updates/5404b92de4b0f2ed1e146373

OpenStudy (mathmath333):

1.)for maximum value of sine function u have to make it 90 degree so try to first make \[120+Q=90\] first which will be when \[q=-30\] then also put \[q=-30\] in the second sine function \[120-Q\] . calculate the result . 2.)now make \[120-Q=90\] which will come when u take \[Q=30\] and also put the value \[Q=30\] in first sine function \[120+Q\] calculate the result. u will find both the result to be same and hence will be the maximum value